# PinRec: Outcome-Conditioned, Multi-Token Generative Retrieval for Industry-Scale Recommendation Systems

### v1

@@ 1. Introduction
Traditionally, retrieval methods employ “two-tower” models. The query tower computes query embeddings, while the item tower computes item embeddings. These models retrieve items with the highest similarity scores given a query embedding. As an alternative to two-tower models, recent work has proposed generative retrieval. Generative retrieval models leverage a sequence of tokens using generative models, such as transformers (Vaswani et al., 2017), and then use those tokens to retrieve items. Generative retrieval has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (Kang and McAuley, 2018; Li et al., 2023; Yang et al., 2024). Nevertheless, to our knowledge, generative retrieval has not been attempted as successfully for large-scale industrial applications yet.
@@
To our knowledge, there has been no large scale generative retrieval model that has addressed all of the aforementioned challenges. Most methods developed within academic settings (Kang and McAuley, 2018; Hou et al., 2022; Li et al., 2023) have not proven effective in meeting the latency and cost requirements necessary to handle web-scale traffic. While there are a few recent studies to apply generative retrieval to production, they focus on optimizing one objective (Zhai et al., 2024) and do not demonstrate optimal optimization for different target outcomes.
@@ 2.1. Sequential Recommendation Systems
Sequential recommendation systems operate on a sequence of items with which a user has interacted (Kang and McAuley, 2018). Earlier work in academia has used recurrent neural networks (RNN) (Hidasi et al., 2016), but later work has often adopted transformers (Kang and McAuley, 2018; Sun et al., 2019). Recent studies have harnessed the extensive knowledge of pre-trained language models (Hou et al., 2022; Li et al., 2023).
@@ 5.2. Offline Results
In this section, we compare PinRec with an existing sequence-based user modeling representation, Pinnerformer (Pancha et al., 2022), in terms of offline evaluation metrics. Further, we quantify the impacts of our design on the performance and diversity of recommendations.
@@ 5.2.1. Comparison with Prior Method
Table 1. Comparison of PinnerFormer and PinRec variants (unordered recall @ 10) across major surfaces at Pinterest.
@@
PinnerFormer 0.461 0.412 0.257
@@ Appendix A Appendix
A.1. Visualization of Multi-Token Impact
@@
(a) Recent user history manually specified to produce results in Subfigures LABEL:subfig:nails_tattoos_user_non_mt and LABEL:subfig:nails_tattoos_user_mt. In reverse chronological order. Note a mix of nail and tattoo-related queries and engagements.
@@
(a) Recent user history manually specified to produce results in Subfigures 9(b) and 9(c). In reverse chronological order. Note a mix of food and nursery-related queries and engagements.
(b) Results for PinRec-OC candidate generator for user history in Subfigure 9(a). Notably, the model is attending more to recent food-related interactions than earlier nursery-related interactions.
(c) Results for PinRec-{MT, OC} candidate generator for user history in Subfigure 9(a). Includes a mix of Pins related to nurseries and food.

### v2

@@ 1. Introduction
Traditionally, retrieval methods employ “two-tower” models. The query tower computes query embeddings, while the item tower computes item embeddings. These models retrieve items with the highest similarity scores given a query embedding. As an alternative to two-tower models, recent work has proposed generative retrieval. Generative retrieval models leverage a sequence of tokens using generative models, such as transformers (Vaswani et al., 2017), and then use those tokens to retrieve items. Generative retrieval has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (Kang and McAuley, 2018a; Li et al., 2023; Yang et al., 2024). Nevertheless, to our knowledge, generative retrieval has not been attempted as successfully for large-scale industrial applications yet.
@@
To our knowledge, there has been no large scale generative retrieval model that has addressed all of the aforementioned challenges. Most methods developed within academic settings (Kang and McAuley, 2018a; Hou et al., 2022; Li et al., 2023) have not proven effective in meeting the latency and cost requirements necessary to handle web-scale traffic. While there are a few recent studies to apply generative retrieval to production, they focus on optimizing one objective (Zhai et al., 2024) and do not demonstrate optimal optimization for different target outcomes.
@@ 2.1. Sequential Recommendation Systems
Sequential recommendation systems operate on a sequence of items with which a user has interacted (Kang and McAuley, 2018a). Earlier work in academia has used recurrent neural networks (RNN) (Hidasi et al., 2016), but later work has often adopted transformers (Kang and McAuley, 2018a; Sun et al., 2019). Recent studies have harnessed the extensive knowledge of pre-trained language models (Hou et al., 2022; Li et al., 2023).
@@ 5.2. Offline Results
In this section, we compare PinRec with an existing sequence-based user modeling representations, Pinnerformer (Pancha et al., 2022), SASRec (Kang and McAuley, 2018b), and TIGER (Rajput et al., 2023), in terms of offline evaluation metrics. Further, we quantify the impacts of our design on the performance and diversity of recommendations.
@@ 5.2.1. Comparison with Prior Method
Table 1. Comparison of baselines and PinRec variants (unordered recall @ 10) across major surfaces at Pinterest.
@@
TIGER (Rajput et al., 2023) 0.208 0.230 0.090
SASRec (Kang and McAuley, 2018b) 0.382 0.426 0.142
PinnerFormer (Pancha et al., 2022) 0.461 0.412 0.257
@@ Appendix A Appendix
A.1. Additional Analysis
In this section, we perform additional experiments to quantitatively and qualitatively show that the design decisions made for PinRec are justified and that it performs across the state-of-the-art architectures. Additionally, we provide more online analyses of the performance across different user cohorts and tail items.
A.1.1. HSTU Architecture
Though PinRec relies primarily on the well-known transformer architecture, we note that our technique is completely architecture-agnostic. Specifically, outcome conditioning and multi-token prediction can be applied regardless of the specific architecture (e.g., transformer blocks) or algorithms used (e.g., attention), allowing them to be generic for any future use case of generative retrieval despite architectural or algorithmic improvements. To demonstrate this, we reproduce significant gains over baselines using the HSTU architecture (Zhai et al., 2024), as opposed to the multi-head attention-based causal transformer. Note that we do not leverage the ranking task or any sequence sampling explored in (Zhai et al., 2024) as it is not relevant for our specific use case. For simplicity, we reproduce results on the outcome conditioned variant of PinRec only, as we believe similar results can be extrapolated to other variants.
The results are shown in Table 4, which clearly shows that HSTU is roughly 3% worse on HomeFeed (effectively negating the improvement with respect to the unconditioned transformer-based PinRec) and competitive on the related pins surface. However, there is a substantial drop in performance on the search surface, which is a critical component of the user experience at Pinterest. We believe that this may be related to specific architectural elements of HSTU.
Table 4. Comparison of HSTU and transformer architectures for PinRec, where PinRec-OC leverages the transformer as presented in the main text.
Model Homefeed Related Pins Search
PinRec-OC 0.625 0.537 0.352
HSTU (OC) 0.596 0.539 0.179
Though we observe a sharp drop in the search surface, HSTU performs roughly on the same order of magnitude as the traditional transformer architecture, demonstrating that our ideas perform across different architectures. This validates that our ideas are robust to architectures across the board.
A.1.2. Ablating Negative Sampling
Following from Pancha et al. (2022), we perform ablations on negative sampling, as it is a critical component of the chosen sampled softmax loss function. Though we ultimately end up with similar conclusions as Pancha et al. (2022), we show these findings for maximum clarity regarding our design choices.
Random Negatives
Though in-batch negatives are a critical component of our negative corpus, we construct a random negative corpus consisting of a random set of possible items. Given that our retrieval problem is ultimately across a random selection of pins, we wish for PinRec to show retrieval capability across this corpus. We perform several ablations with random negatives, with in-batch negatives constant at a non-zero value. For these analyses, we focus on Homefeed (the most critical surface) and present results aggregated across all surfaces as well.
We show results in Table 5, where there are moderate gains in unordered recall if there is a non-zero number of in-batch negatives present. Specifically, on the most critical surface of Homefeed, we see notable improvement over the baseline in terms of unordered recall.
Table 5. Percentage lift in unordered recall using 16K random negatives, relative to using only 8K random negatives.
Surface % Lift in Recall @ 10
Homefeed +2.5%
All +1.3%
To summarize, while random negatives are certainly an important piece of the puzzle, they cannot substitute in-batch negatives. When used in conjunction, scaling the number of negatives yields notable improvement, especially on Homefeed.
In-Batch Negatives
As shown in the previous subsection, disabling in-batch negatives entirely leads to significantly poorer metrics, showing their importance. Can we improve performance offline by scaling the batch size and consequently, the number of in-batch negatives? To examine this, we fix the number of random negatives and ablate the number of in-batch negatives through the batch size.
Table 6. Percentage lift in unordered recall @ 10 through scaling the number of in-batch negatives (relative to baseline of 1x) with random negatives fixed.
# In-Batch Negs. % Lift (HF) % Lift (all)
1.5x +4.2% +3.3%
2x +8.4% +5.4%
12x +18.3% +12.1%
We show results in Table 6, which demonstrates a clear trend that shows that a greater number of in-batch negatives and batch size leads to better performance, especially for the Homefeed surface. These are similar in directional improvements to random negatives, but we observe greater improvements with in-batch negatives, which are likely more relevant as negative examples.
A.1.3. Ablating Parameter Scaling
One of the recent explorations within recommendation systems has been scaling to improve performance, in an attempt to replicate the success of these efforts for language modeling. To explore this hypothesis, we scale the number of parameters within PinRec across two fronts: the “vocabulary” and the model itself. While model scaling is clear and involves scaling up the embedding size, the number of transformer layers, and so on, scaling the vocabulary for an embedding-based generative retrieval technique involves improving the input representations. Specifically, as mentioned in the main text, we implement “ID” embeddings for pins, that effectively serve as additional representation for the “pin vocabulary”.
Model Parameters
As presented in the main text, PinRec (all variants) have roughly 100 million transformer parameters, corresponding to the base form or size of GPT-2. To ablate the effect of the size of the transformer, we attempt to follow the different scales employed by GPT-2 (base, medium, large, XL) as they have shown success in terms of the factors by which the embedding size, number of layers, and number of heads have scaled. Note that due to computational constraints, we reduce the scaling in terms of the number of layers, which increases GPU memory footprint significantly. Note that we present lift in offline unordered recall @ 10 for the outcome conditioned variant of PinRec below, but we observe similar trends for the unconditioned model as well.
Table 7. Percentage lift in unordered recall through expanding the transformer at various GPT-2 scales relative to base ( 100 million parameters).
Model % Lift in Recall @ 10
Medium -0.8%
Large -0.9%
XL +2.2%
Though we observe some gains in Table 7 from scaling the number of transformer parameters to around 1 billion (XL-sized), these become intractable to serve in real time, requiring offline batch inference or significantly higher cost and more optimization. For the remaining sizes, we observe insignificant lifts (note that less than 1% in lift magnitude can be often explained through noise). Hence, we opt to not use larger sized transformers.
Vocabulary Parameters
As presented in the main text, the ID embeddings consist of roughly 10 billion parameters. These are a core component of the model since they provide additional representational power for the most important entity type at Pinterest: Pins. To ablate the importance of these ID embeddings, we examine the performance in offline recall across all surfaces with and without ID embeddings. For this analysis, we focus on Homefeed (the most critical surface) and present results aggregated across all surfaces as well in Table 8.
Table 8. Percentage lift through adding ID embeddings (10 billion parameters) over no ID embedding-based vocabulary.
Surface % Lift in Recall @ 10
Homefeed +14.0%
All +14.4%
Evidently, the improvement in recall is substantial and indicates that the scaling of vocabulary to learn an embedding for each unique ID (roughly, as it is hash-based) rather than only content-based features is beneficial.
A.1.4. Tail User and Item Analysis
One of the potential drawbacks of outcome conditioning would be for “tail” or “fresh” users, i.e., those which have lesser activity or less propensity for high-effort engagement. Critically, our method intends to optimize for outcomes across all cohorts of users, including such users. In the same vein, we wish to optimize for “tail” items, which may receive a lesser number of impressions, in an effort to distribute content to improve corpus coverage.
User Cohort Grid Clicks Outbound CTs
UC OC UC OC
New +2.60% +2.62% +3.98% +5.78%
Resurrected +1.56% +1.86% +1.98% +1.72%
(a) Online lift (%) for grid clicks and outbound clickthroughers for new and resurrected users.
Metric UC OC
Corpus Coverage (unique pins) +6.2% +17.3%
1-Imp Pin Rec. (tail items) +5.8% +7.7%
(b) Relative lift (vs. 2-tower model) for tail item and corpus coverage metrics.
Table 9. Online experimental results for different user cohorts and tail item metrics, across PinRec-UC and PinRec-OC.
In terms of new users, PinRec-UC/OC lift grid clicks and outbound clickthroughers (outbound CTs) in 9(a). Critically, the lift for outcome conditioning-based modeling is either neutral or positive with respect to both control and the unconditioned variant of PinRec, demonstrating that outcome conditioning does not suffer on such cohorts of users. In fact, across new users, OC performs better on outbound CTs.
To examine performance for tail items, we compare to a production 2-tower model in an online setting in Table 9. On corpus coverage, adding PinRec-UC/OC surfaces 6.2%/17.3% more unique pins relative to the 2-tower. For long-tail items with 1-impression, PinRec-UC/OC achieve 5.8%/7.7% absolute lifts respectively, relative to the aforementioned 2-tower model. Evidently, across the board, PinRec surfaces more of the corpus and surfaces more tail items compared to control, with outcome conditioning showing even more significant lifts over control in both regards.
We believe the evidence suggests that our generative retrieval solution is equal to or better across inactive/new users and tail items relative to existing control, and perhaps more importantly, it shows that outcome conditioning can help in boosting performance in these criteria, which is one of its goals.
A.1.5. Serving Efficiency
One of the critical aspects of our proposed generative retrieval technique is an effort to maintain serving efficiency. With our custom infrastructure built from off-the-shelf serving technologies such as NVidia Triton Inference Server and optimization techniques like torch.compile, we achieve a system that can be served in real-time. In this section, we delve deeper into the memory and time requirements, comparing our method with prior methods deployed in production.
On the memory front, all variants of PinRec (unconditioned, outcome-conditioned with or without multi-token generation) use about 1.6GiB of GPU memory including KVCache and CUDA graphs, without any significant memory usage increase for multi-token prediction or outcome conditioning. Excluding the KVCache and CUDA graphs, we use roughly the same amount of memory (2̃50 MiB) relative to other techniques (i.e., two-tower models).
In terms of latency, at 80 QPS, the online and offline latency for PinRec-OC/OC, MT are p50 40ms/p90 65ms (6 steps). On the other hand, across only 3 steps for PinRec-UC, we observe a latency of p50 23ms and p90 46ms (note that this produces significantly less embeddings and less compressed embeddings than either of the above techniques, though it is lower latency). Though the latency is roughly 3-4x higher (p50) than a traditional two-tower solution, we note that parallelizing it with other RPCs results in ¡1% added end-to-end increase in latency.
A.2. Visualization of Multi-Token Impact
@@
(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of nail and tattoo-related queries and engagements.
@@
(a) Recent user history manually specified to produce results in subfigures. In reverse chronological order. Note a mix of food and nursery-related queries and engagements.
(b) Results for PinRec-OC candidate generator for given user history. Notably, the model is attending more to recent food-related interactions than earlier nursery-related interactions.
(c) Results for PinRec-{MT, OC} candidate generator for given user history. Includes a mix of Pins related to nurseries and food.